# MC-MoE: Mixture Compressor for Mixture-of-Experts LLMs Gains More

### v1

@@ 4.2 Experiment on Online Dynamic Pruning
Ablation of Tokens Protection. As shown Fig. 8, when we select 2% crucial tokens to be protected, the PPL drops from 6.46 to 6.24, with activated experts’ parameters decreasing only from 15.1% to 14.8%. Moreover, as we gradually increase the ratio, the performance remains relatively stable, while the compression ratio exhibits a nearly linear decline. Thus, we conclude that protecting just 2% of the important tokens can significantly enhance the compressed performance of MoE-LLMs with almost no impact on efficiency. Furthermore, we try to prune all experts associated with the less important tokens, as illustrated in Fig. 8. When removing the experts to the 2% of tokens, the overall compression ratio reached 15.8%, while the PPL improved to 6.35, yielding performance enhancements in both efficiency and accuracy compared to weight-only pruning. However, we observed that the performance curves of all experts’ masking exhibited exponential growth, indicating that directly skipping experts results in a significant accuracy loss. By employing a protection mechanism for only 2% of the experts, we can maintain the accuracy of MoE-LLMs without compromising efficiency.
Memory Saving and Inference Efficiency. Tab. 4 details the memory compression and speed tests of the proposed MC-MoE. The 16-bit Mixtral $8\times 7$b model requires two A100-80G GPUs, while the Mixtral $8\times 22$b model needs four. We utilize the HQQ (Badri & Shaji, 2024) tool to save quantized weights and handle dequantization. To saving the binary weight, we design a bit-change transformation (see Appendix A.2). After applying PMQ mixed-precision compression, the Mixtral $8\times 7$b model can be compressed to a size ranging from 10.82 to 16.65 GB. During dynamic inference, ODP reduces activation parameters by about 15%, with average accuracy decreasing by less than 1%. At 2.05-bit, the average activation parameter per token is only 3.23 GB, resulting in a 1.80×1.80$\times$1.80 × increase in inference speed and an evaluation accuracy of 62.68%. Tab. 4 also compares the performance of the LLaMA series dense models. The MC-compressed 2.54-bit Mixtral $8\times 7$b model outperforms the 26.03 GB 16-bit LLaMA2-13b model, with a total parameter size of 16.65 GB and activation parameters of 3.69 GB. We have also extended the compression experiments to the Mixtral $8\times 22$b model. MC-MoE shows higher overall performance compared to mainstream dense models, demonstrating the potential for compressing MoE-LLMs without additional training.
@@ 5 Conclusion
MoE represents a promising framework of sparse models for natural language understanding through scaling up the model capacity. However, the memory demands and redundancy among experts pose significant challenges for their practical implementation. In this work, we propose MC-MoE, a mixture compression strategy based on the imbalance of significance among experts. This method co-designs the Pre-Loading Mixed-Precision Quantization (PMQ) and Online Dynamic Pruning (ODP) approach, allowing MoE models to be compressed to an ultra-low bit-width while maintaining exceptional memory and parameter efficiency, as well as knowledgeable performance. And our mixed-precision strategy is orthogonal to various quantization techniques. Comprehensive experiments validate the effectiveness of our mixture compression, revealing that highly compressed MoE-LLMs can even outperform equal-size full-precision dense LLMs, thereby improving the feasibility of MoE compression. Future work will focus on adapting this strategy for multimodal applications and optimizing it for specific hardware platforms.
@@ A.1 More Quantized Results of PMQ
Table 5: Performance of quantized Mixtral $8\times 7$b on eight zero-shot benchmarks. ”HellaS.” is the short format of ”HellaSwag” and ”Wino.” denotes ”Winogrande”.
@@
This paper presents MC-MoE, which explores static compression strategies and dynamic pruning method for MoE-LLMs in the ultra-low bit-width range, with selected static bit-width of 1-bit, 2-bit, and 3-bit. We observe that both 2-bit and 3-bit can be addressed using conventional linear quantizers, a method commonly utilized in most studies (Frantar et al., 2022; Shao et al., 2023; Huang et al., 2024c; Lin et al., 2024). In contrast, the quantization of 1-bit weights involves totally different calculations; we first provide the binarization formula for the weights:
@@ A.3 Results of Different Quantization Techniques
Table 8: Performance of quantized Mixtral $8\times 7$b on eight zero-shot benchmarks on GPTQ (Frantar et al., 2022) and Omniquant (Shao et al., 2023). w denotes ”with”.
@@
A.4 Detailed Results on Bit-Width Allocation
@@
A.5 Ablation Analysis on Hyper-parameters of Expert Significance Weight

### v2

@@ 4.2 Experiment on Online Dynamic Pruning
Ablation of Tokens Protection. As shown Fig. 8, when we select 2% crucial tokens to be protected, the PPL drops from 6.46 to 6.24, with activated experts’ parameters decreasing only from 15.1% to 14.8%. Moreover, as we gradually increase the ratio, the performance remains relatively stable, while the compression ratio exhibits a nearly linear decline. Thus, we conclude that protecting just 2% of the important tokens can significantly enhance the compressed performance of MoE-LLMs with almost no impact on efficiency. Furthermore, we try to prune all experts associated with the less important tokens, as illustrated in Fig. 8. When removing the experts to the 2% of tokens, the overall compression ratio reached 15.8%, while the PPL improved to 6.35, yielding performance enhancements in both efficiency and accuracy compared to weight-only pruning. However, we observed that the performance curves of all experts’ masking exhibited exponential growth, indicating that directly skipping experts results in a significant accuracy loss. By employing a protection mechanism for only 2% of the experts, we can maintain the accuracy of MoE-LLMs without compromising efficiency. We also provide the detailed ablation on pruning threshold in Appendix A.8.
Memory Saving and Inference Efficiency. Tab. 4 details the memory compression, speed tests, and average results (Gao et al., 2013) (LM-Eval) of the proposed MC. The 16-bit Mixtral $8\times 7$b model requires two A100-80G GPUs, while the Mixtral $8\times 22$b model needs four. We utilize the HQQ (Badri & Shaji, 2024) tool to save quantized weights and handle dequantization. To saving the binary weight, we design a bit-change transformation (see Appendix A.2). After applying PMQ, the Mixtral $8\times 7$b model can be compressed to a memory from 10.82 to 16.65 GB. During dynamic inference, ODP reduces activation parameters by about 15%, with average accuracy decreasing by less than 1%. At 2.05-bit, the average activation parameter per token is only 3.23 GB, resulting in a 1.80×1.80$\times$1.80 × increase in inference speed and an evaluation accuracy of 62.68%. Tab. 4 also compares the performance of the LLaMA series dense models. The MC compressed 2.54-bit Mixtral $8\times 7$b model outperforms the 26.03 GB 16-bit LLaMA2-13b model, with a total parameter size of 16.65 GB and activation parameters of 3.69 GB. We have also extended the compression experiments to the Mixtral $8\times 22$b model. MC shows higher overall performance compared to mainstream dense models, without any training of original model. (more real speedup results are shown in Appendix A.9).
@@ 5 Conclusion
MoE represents a promising framework of sparse models for natural language understanding through scaling up the model capacity. However, the memory demands and redundancy among experts pose significant challenges for their practical implementation. In this work, we propose MC, a mixture compression strategy based on the imbalance of significance among experts. This method co-designs the Pre-Loading Mixed-Precision Quantization (PMQ) and Online Dynamic Pruning (ODP) approach, allowing MoE models to be compressed to an ultra-low bit-width while maintaining exceptional memory and parameter efficiency, as well as knowledgeable performance. And our mixed-precision strategy is orthogonal to various quantization techniques. Comprehensive experiments validate the effectiveness of our mixture compression, revealing that highly compressed MoE-LLMs can even outperform equal-size full-precision dense LLMs, thereby improving the feasibility of MoE compression. Future work will focus on adapting this strategy for multimodal applications and optimizing it for specific hardware platforms.
Acknowledgments
This work has been supported in part by Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923), Theme-based Research (Grant No. T45-701/22-R), the Innovation and Technology Fund (Mainland-Hong Kong Joint Funding Scheme, MHP/053/21), and the Shenzhen-Hong Kong-Macau Technology Research Program (SGDX20210823103537034). This research is also supported in part by National Key R&D Program of China (2022ZD0115502), National Natural Science Foundation of China (NO. 62461160308, U23B2010), ”Pioneer” and ”Leading Goose” R&D Program of Zhejiang (No. 2024C01161).
@@ A.1 More Quantized Results of PMQ
Table 5: Performance of quantized Mixtral $8\times 7$b on eight zero-shot benchmarks. “HellaS.” is the short format of “HellaSwag” and “Wino.” denotes “Winogrande”.
@@
This paper presents MC, which explores static compression strategies and dynamic pruning method for MoE-LLMs in the ultra-low bit-width range, with selected static bit-width of 1-bit, 2-bit, and 3-bit. We observe that both 2-bit and 3-bit can be addressed using conventional linear quantizers, a method commonly utilized in most studies (Frantar et al., 2022; Shao et al., 2023; Huang et al., 2024c; Lin et al., 2024). In contrast, the quantization of 1-bit weights involves totally different calculations; we first provide the binarization formula for the weights:
@@ A.3 Results of Different Quantization Techniques
Figure 9: Needle in a Haystack evaluation. Green squares indicates a high retrieval success rate, the Y-axis represents the distance to the retrieved target.
Table 8: Performance of quantized Mixtral $8\times 7$b on eight zero-shot benchmarks on GPTQ (Frantar et al., 2022) and Omniquant (Shao et al., 2023). w denotes “with”.
@@
A.4 Quantization Results on Challenging Benchmarks
In this section, we expand our mixed-precision benchmarks on more challenging datasets in Tab. 9, considering the importance of performance testing on more complex long text or reasoning tasks (Cobbe et al., 2021; Bai et al., 2023; Chen et al., 2021). We have observed that in challenging tasks like GSM8K, HumanEval, and long-context Needle-in-a-haystack, the performance drop of model compression becomes more pronounced. This phenomenon holds true in other MoE LLM compression methods (Frantar et al., 2022; Huang et al., 2024c; Shao et al., 2023; Lu et al., 2024) as well. However, our PMQ method, compared to the latest method like BSP (Li et al., 2024) and HAWQ (Dong et al., 2020) with Hessian-based approaches for MoE LLM, is still able to maintain state-of-the-art performance. Fig. 9 shows the NIAH results in different sequences.
Recent studies on quantization performance losses (Jin et al., 2024; Liu et al., 2023a) were also explored, revealing that ARC-C and GSM8K primarily involves inference issues, categorized as chain-of-thought (CoT), while MMLU can be classified as in-context learning (ICL). CoT tasks, due to their intricate reasoning demands, pose significant challenges to various LLM types. Given that many open-source MoE LLMs and dense LLMs do not exhibit strong inference capabilities during pre-training, we anticipate larger performance losses when reducing model bit-width to ultra-low scenarios. The results in Tab. 2 and Tab. 9 also indicate that there is the huge potential for future exploration of MoE LLM compression on complex tasks.
Table 9: Comparison of different mixed-precision quantization methods on challenging benchmarks. NIAH denotes the task in Needle-in-a-haystack, which is a more challenging task for evaluating long-context ability.
[…]
16.00 58.30 59.15 100.00
Uniform 3.00 38.13 29.88 98.48
Uniform 2.00 0.00 0.00 0.00
BSP 2.54 4.25 3.21 42.21
Hessian 2.54 33.59 25.49 100.00
Hessian 2.05 17.24 7.84 93.45
PMQ 2.54 37.67 29.34 100.00
PMQ+ODP 2.54 35.25 27.58 100.00
[…]
PMQ+ODP 2.05 18.04 10.02 99.26
A.5 Detailed Results on Bit-Width Allocation
@@
A.6 Ablation Analysis on Hyper-parameters of Expert Significance Weight
@@
A.7 Comparison of Different Token-dependent Pruning Metric
Regarding the dynamic pruning of experts, we note that most existing pruning methods for LLMs or other neural networks focus on static weight pruning (Sun et al., 2023; Zhang et al., 2023), and cannot dynamically prune experts during inference based on tokens. Dynamic pruning during inference remains under-explored, with only one recent post-training MoE LLM dynamic pruning work (Lu et al., 2024) proposing a gating-score-based strategy for dynamic pruning. This work has already compared with Wanda (Sun et al., 2023) method, a highly effective static pruning method, and concluded that static pruning methods result in significant performance degradation when applied to dynamic MoE LLM experts. We incorporated additional metrics for dynamic expert pruning to expand the scope of our experiments. Specifically, we perform token-dependent expert pruning on token kurtosis, token var, and token mean, where 30% of tokens will be pruned from $top\mbox{-}2$ to $top\mbox{-}1$ (Tab. 11).
Table 11: Comparison of different token-dependent dynamic expert pruning strategies on Mixtral $8\times 7$b. Avg.CP denotes the average compressed parameters ratio for each token. NIAH denotes the task in Needle-in-a-haystack, which is a more challenging task for evaluating long-context ability.
[…]
Token kurtosis 0.3 15.62% 7.16 57.22 14.05 6.54 93.16
Token variance 0.3 15.62% 6.69 60.02 17.33 7.92 95.37
Token mean 0.3 15.62% 6.82 59.27 17.76 6.02 95.65
ODP - 14.88% 6.22 63.25 18.04 10.02 99.26
A.8 Ablation of Dynamic Expert Pruning Threshold
We follow the setting from recent dynamic MoE pruning work (Lu et al., 2024), selecting it as the median value of $\frac{w_{1}}{w_{0}}$, which also theoretically and empirically demonstrates that this choice of threshold is a comprehensive optimal setting. In this section, we provide a more comprehensive ablation on threshold $\mu$ in Eq. 5. As demonstrated in Table 12, utilizing a manual threshold of 0.4, the PPL performance stands at 6.29 with a mere 12.00% of experts pruned. In contrast, our proposed method, referred to as ODP, achieves a PPL of 6.22 and prunes 14.88% of the experts. This not only showcases superior accuracy but also highlights enhanced efficiency.
Table 12: Ablation of different threshold hyperparameter.
$\mu(w_{1}/w_{0})$ PPL (WikiText2$\downarrow$ Avg. Pruning Params.
0.4 6.29 12.00%
0.5 6.49 16.51%
0.6 6.64 19.25%
0.7 6.89 22.43%
Median 6.48 15.18%
ODP(Median + Protection) 6.22 14.88%
A.9 Computation Analysis of Online Dynamic Pruning
During the ODP phase, compared to the significant reduction in the number of tokens and experts leading to large-scale matrix multiplications, the computational cost of token importance calculation can be negligible. Specifically, in Mixtral $8\times 7$b, where the typical input token matrix size is $R^{n\times m}$, the token importance calculation involves three steps: summing attention weights, computing the ${\ell_{1}}$ norm, and performing $top\mbox{-}k$ calculations. The overall floating-point operations per second (FLOPs) calculation amounts to $n^{2}+n+mn+nlogn$. In the ODP inference phase, after dynamic pruning, an average of 15% of tokens in a MoE layer will reduce an experts inference (see Tab. 12). The FLOPs for these 15% of tokens within an expert (an expert with 3 linear layers, the size is $R^{m\times m_{1}}$, $R^{m_{1}\times m_{1}}$, $R^{m_{1}\times m}$) are $0.15n*\times(m\times m_{1}\times 2+m_{1}^{2}\times 2+m_{1}\times m\times 2)$, where $m_{1}$ is typically much larger than n and m in Mixtral $8\times 7$b. Therefore, the computational cost of importance calculation is usually low. As demonstrated in Tab. 4, when PMQ is combined with ODP, it further enhances computational efficiency. This indicates that the efficiency gain from experts’ dynamic pruning outweighs the computational cost of token importance calculation.
Table 13: End-to-end latency comparison between FP16 and MC on Mixtral $8\times 7$b under different [batch, input token length]. Each cell is the latency for one token generation speed (second).
Hardware [1,512] [1,1024] [1,2048] [1,4096] [8,2048] [8,4096] [16,2048] [16,4096]
FP16 2×2$\times$2 ×A100 0.029 0.038 0.043 0.057 0.009 0.011 0.007 0.010
MC 2.54-bit 1×1$\times$1 ×A100 0.015 0.018 0.019 0.025 0.004 0.005 0.004 0.004
Speedup(%) - 48.3 52.7 56.2 56.1 55.4 54.3 47.6 60.1
Table 14: Latency comparison of MoE and dense LLM under different hardware platform.
Model Hardware Loading Memory Peak GPU Memory LM-Eval%$\uparrow$ Token/s
Mixtral 8$\times$7b 2×2$\times$2 ×A100 96.8 GB 112.6 GB 71.29 23
Mixtral 8$\times$7b 1×1$\times$1 ×3090 OOM OOM - -
LLaMA2-13b 1×1$\times$1 ×A100 26.0 GB 33.4 GB 65.19 46
LLaMA2-13b 1×1$\times$1 ×3090 OOM OOM - -
Mixtral 8$\times$7b MC 2.54-bit 1×1$\times$1 ×A100 16.2 GB 20.7 GB 66.94 38
[…]
In Tab.13 and Tab.14, we present the actual speed enhancements of deployment achieved by our MC method on various hardware platforms. The speed enhancements in MC, as detailed in Tab. 13, originate from static compression during the PMQ phase and adaptations of the CUDA kernel (based on HQQ) along with ODP. Across varying batch sizes and input sequence lengths, our speedup ranges from 40% to 60%. The performance boost from model weight compression remains consistent regardless of input sequence length and batch size. However, with a fixed batch size, we notice a more pronounced speed advantage for our MC-MoE as the sequence length increases, attributed to the increased efficiency demonstrated by ODP. As batch size increases, both the FP16 models and compressed models experience an overall increase in throughput, leading to accelerated average token generation speeds. In Tab. 14, with MC-MoE on the RTX 3090 GPU, extreme compression allows for an average speed of 52 token/s, which is very cost-effective. In this scenario, compressed MoE LLM outperforms dense LLM in memory, accuracy, and speed.